# Reinforcement Inference: Leveraging Uncertainty for Self-Correcting Language Model Reasoning

### v1

@@ X Conclusion
We introduced Reinforcement Inference, an inference-time approach that improves an LLM’s answers by leveraging its own uncertainty signals. Through experiments on 12,032 challenging multiple-choice questions from MMLU-Pro across 14 subjects, we demonstrated that when DeepSeek-v3.2 exhibited high predictive entropy—a proxy for hesitation—that uncertainty often signaled an incorrect first-pass answer, and that a second, more deliberate attempt could correct a large fraction of these errors without any retraining.
This work highlights several key insights. First, large language models like DeepSeek do have measurable internal confidence; in our tests at scale, wrong answers were usually accompanied by higher entropy and lower max probability than correct answers. This underscores the importance of developing techniques to read and use those internal signals for more reliable AI behavior. Second, a simple second-pass prompt (essentially telling the model, ”you were unsure, so be careful now”) can unlock better reasoning in the model. The success of this prompt shows how far prompt-based control can go in steering model cognition–in this case, nudging the model from a fast guess toward a more deliberative mode. Third, our method achieved substantial gains at scale: overall accuracy rose from 60.72% to 84.03% (+23.31%) while re-asking 61.06% of questions. A uniform re-asking ablation (UR) in which we re-asked 100% of questions yielded only +0.10% additional accuracy over targeted re-asking (TR), indicating that uncertainty-based triggering is the key–not the volume of re-asking–and that uncertainty is central to addressing hallucination when the LLM is confused. Nearly identical results with 39% less computation demonstrate both the efficiency and the correctness of our design. Crucially, all of this was done at inference time, preserving the model’s weights. This has practical implications: even when one cannot retrain a model (due to cost, or if using a fixed API model), it may still be possible to improve its reliability through intelligent post-processing or on-the-fly prompting of its outputs.
Reinforcement Inference contributes to the evolving toolbox for aligning and controlling LLM outputs [34]. It resonates with the idea of making models not just give an answer, but also reflect on how they arrived at it and whether they are satisfied with it. By operationalizing a form of self-reflection triggered by uncertainty, we move a step closer to LLMs that can know what they know–and know what they don’t. We hope this work inspires further research into uncertainty-aware AI and meta-cognitive prompting techniques, which together could make AI systems not only more accurate, but also more self-aware and trustworthy in their reasoning.
More broadly, our method is tightly coupled to the autoregressive nature of modern decoder-based models: because answers (and more generally, sequences) are generated token by token from a learned conditional distribution, entropy and related confidence measures arise naturally as first-class control signals during generation. Reinforcement Inference exploits this coupling by turning distributional uncertainty into an adaptive inference policy, allocating an additional pass only when the model’s own posterior indicates hesitation, thereby converting internal ambiguity into measurable and actionable compute. The same uncertainty-triggered refinement principle naturally extends beyond discrete answer selection to sequence-generation settings, including vision–language–action (VLA) models and robotic control. In such systems, spatial or action entropy can be interpreted as a proxy for execution uncertainty: high entropy corresponds to diffuse, unstable action proposals, whereas low entropy indicates confident, convergent trajectories. This perspective closely parallels human motor behavior, where early movements are exploratory and imprecise, progressively concentrating into precise actions as the hand approaches the target, motivating uncertainty-aware action refinement as a foundation for improving reliability in autoregressive VLA generation.
Stepping back, the broader message is that uncertainty is not merely a diagnostic; it is a control signal. Treating entropy as a first-class input to an adaptive inference policy provides a principled way to allocate compute, probe a model’s latent capability horizon, and improve reliability without changing model weights.

### v3

@@ X Conclusion
We introduced Reinforcement Inference, an inference-time method that leverages a model’s uncertainty to selectively trigger a second, more deliberate attempt.
On 12,032 MMLU-Pro multiple-choice questions across 14 subjects, DeepSeek-v3.2 improves from 60.72% to 84.03% (+23.31%) while re-asking 61.06% of questions; re-asking 100% yields only +0.10% additional accuracy over targeted re-asking (TR), indicating that uncertainty-based triggering captures most of the gain with substantially less compute.
More broadly, our framework is naturally aligned with the autoregressive nature of modern decoder-based models: because outputs are generated token by token from a learned conditional distribution, entropy and related confidence measures arise directly during generation and can be used as control signals. Reinforcement Inference exploits this coupling by turning distributional uncertainty into an adaptive inference policy—allocating extra compute only when the model hesitates—to improve reliability without changing model weights. This perspective complements broader efforts to align and control LLM behavior and motivates further work on uncertainty-aware inference and training.